# Towards Linear Spanners in All Temporal Cliques

### v1

@@ 4 Structure of Bi-Clique Instances
Using Theorem 3.1, from now on we only consider bi-cliques. Our first structural insight for this type of instance was already used by Casteigts, Peters, and Schoeters [6] on special reduced instances of bi-cliques. Remember, that we call an instance extremally matched if the set of all earliest edges $\operatorname{\pi^{\text{--}}}$ and the set of all latest edges $\operatorname{\pi^{+}}$ each form a perfect matching. We give a general reduction rule which reduces the graph in case the instance is not extremally matched. This structural property, that we can assume the existence of such a perfect matching or reduce the instance, is useful for almost all further constructions. For example, it enables us to reduce the instance size, when the number of vertices per side is not equal. Additionally, we can think of the sets $A$ and $B$ of the bi-clique as interchangeable. For example, consider any set of edges that includes temporal paths from some set $A^{\prime}\subseteq A$ to all vertices in $A$. We can extend these paths to reach all vertices in $B$ by including the set of all latest edges, because it is a perfect matching between $A$ and $B$.
@@
Note that this corresponds to the definition in [6], but transferred to the case of bi-spanners, where all vertices in $A$ only need to reach vertices in $B$. Concretely, this means that vertex $a$ can reach vertex $a^{\prime}$ via the earliest edge of $a^{\prime}$ on a temporal path of length 2. Therefore, if this configuration occurs in the bi-clique, we can remove $a$ and include the edges $\mathopen{}\mathclose{{}\left\{{a},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}}\right\}$ and $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}}\right\}$ into our bi-spanner. In the final solution for the whole bi-clique, vertex $a$ then has a temporal path to every vertex $b\in B$ via vertex $a^{\prime}$. Similarly, if a vertex $b\in B$ is dismountable, every temporal path that ends in a different vertex can be extended to be a temporal path ending in vertex $b$, and we can remove $b$ from the bi-clique.
@@
Since from now on, both partitions of $D$ have the same size after removing dismountable vertices, we write $n\coloneqq|A|=|B|$ and call $n$ the size of the instance. This justifies also that the definition of $\operatorname{\mathcal{D}}$ assumes equal partition sizes and gives the following corollary.
@@ 5 Pivotable Bi-Cliques
However, there are instances without this structure [6]. Thus, we study an even more widely applicable structure, based on a what we call a pivot-edge. We exploit pivot-edges to apply the divide-and-conquer paradigm to find linear-size temporal spanners. To the best of our knowledge, this paradigm has not yet been applied to this problem. As we see later, the absence of a pivot-edge yields strong structural properties and it heavily restricts the class of temporal cliques where no linear spanner can be found using current techniques.
5.1 Pivot-Edges
@@
Let $e=\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}\in{A}\otimes{B}$. All vertices in $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ can be organized in a tree $T$ rooted at $a$, s.t. $T$ only has edges with label at least $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$ and for each $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ there are temporal paths $a\leadsto v$ and $b\leadsto v$ in $G[T]$. The analogous result holds for vertices in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$.
The proof of Lemma 5.3 is already known [19] and it relies on constructing a foremost tree for $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ (and its inverted version for $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right))$. It was considered in [16] and lately has been used in [7], where foremost trees are applied to temporal cliques.
We can directly use this idea to temporally connect all $u\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ to all $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ using a linear number of edges.
@@
As $32\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+2\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-3\leq 8n-$ 3, we see that only a linear number of edges is included to split the instance. With this, we can show that if for some edge $e\in{A}\otimes{B}$ the sets $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ and $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ overlap, we can reduce the instance significantly. We name the set ${\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}$ the pivot-set of $e$. If $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq cn$ for $c\in(0,1]$, we call $e$ a $c$-pivot-edge.
Theorem 5.8 (Pivot Edge).
Let $c\in(0,1]$. Let $D=(A,B,\lambda)$ be an extremally matched bi-clique with $n\coloneqq|A|=|B|$. If there is a $c$-pivot-edge in $D$, we can split the instance into two sub-instances with overall number of vertices reduced by $cn$. For each removed vertex, we include at most $\frac{8}{c}$ edges. If applicable recursively, we get a spanner of size $\frac{16}{c}n$.
@@
Let $e\in{A}\otimes{B}$ be such that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq cn$. Note that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|=\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cup\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq\mathopen{}\mathclose{{}\left(2+c}\right)n$. By Theorem 5.6, the sub-instances have sides of overall size at most $2n-\frac{2+c}{2}n=\mathopen{}\mathclose{{}\left(1-\frac{c}{2}}\right)n$. This means that $cn$ vertices have been removed using at most $38n-$ 3 edges overall. For each such vertex, we include $\frac{8n-3}{cn}\leq\frac{8}{c}$ edges.
To recursively apply Theorem 5.6, observe that we can reduce any dismountable vertex until both sub-instances are extremally matched. For each vertex removed in this way, we include $2\leq\frac{8}{c}$ edges. As the number of edges included per removed vertex in both ways does not depend on $n$, we include at most $\frac{8}{c}\mathopen{}\mathclose{{}\left|A\sqcup B}\right|=\frac{16}{c}n$ edges.
@@
5.2 Graph Properties leading to Pivot-Edges
Pivot-edges generalize pivotable vertices and yield a number of new, interesting structural properties in residual instances. In this section, we examine structures that, if present, imply that there is at least one pivot-edge.
We first look at one particular case of pivot-edges, which we call steep edges. Intuitively, an edge is steep, if it belongs to the later edges for one of the incident vertices and to the earlier edges for the other one.
@@ 5.2 Graph Properties leading to Pivot-Edges
While steep edges demonstrate how an intuitively common structure cannot be there if assume there are no pivot-edges. However, in randomly generated instances we expect edges not to be steep. Assume we add edges randomly to the graph in the order of their labels. Then, when adding the edge $\mathopen{}\mathclose{{}\left\{v,w}\right\}$, both $v$ and $w$ should already have a similar number of edges. In fact, one can prove that with high probability, there is no $c$-steep in the graph.
Yet, as every $c$-steep edge is pivotable, we want to analyze the graphs without such edges in greater detail. We first look at more simple properties showing that any node only has a limited number of neighbors for which it is at least the $i$-th and at most the $j$-th neighbor.
@@ 6 Bi-Cliques With Reverted Edges
Now that we have introduced the concept of $c$-pivot-edges and observed some interesting structural properties in bi-cliques which contain no suitable $c$-pivot-edge. The first hope would be that every bi-cliques admits such a $c$-pivot-edge. Unfortunately, this is not the case and we provide a general graph class, the shifted matching graph, in which for every edge $e$, the pivot-set is as small as possible (i.e., $\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)=\{a,b\}$) and none of our reduction rules from Section 4 are applicable.
Still, we present a novel technique, called $e$-reverted edges, to construct small spanners for shifted matching graphs which also provides solutions for a significantly larger class of bi-cliques. We relate the technique to $c$-pivot-edges from Section 5 and argue why $e$-reverted edges solve a class distinct from $c$-pivot-edges, beyond the concrete example of shifted matching graphs. Unfortunately, also this is not applicable to all graph classes, which we later demonstrate with a class that we call product graphs in Section 7.
@@
Figure 4 shows the shifted matching graph for $n=4$ vertices on both sides. In shifted matching graphs, the earliest and latest neighbors form a matching, thus we cannot apply the reduction from Theorem 4.7. Note that for all $i\in\{0,\ldots,{n}-1\}$, we have $\pi^{-}(a_{i})=b_{i}$ and $\pi^{+}(a_{i})=b_{i-1\operatorname{mod}n}$. Additionally, we observe that each edge at time $t$ has the form $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{i+t\operatorname{mod}n}}}\right\}$. Thus, the edges with label $t$ form a perfect matching and $a_{i}$ sees the $b_{j}$ in ascending order of indices, rotated by $i$ places compared to $a_{0}$.
@@
We claim that this graph has no $c$-pivot-edges that are suitable for reduction in the sense of Theorem 5.8. To prove that this method fails on the shifted matching graph, we first characterize the set of vertices which can be reached from a vertex $a_{i}$ before and after a given time. Finally, we prove that there, the intersection of $\operatorname{In}$ and $\operatorname{Out}$ for all edges at $a_{i}$ is trivial.
@@
Using Lemma 6.2, we can conclude that using $c$-pivot-edges in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ leads to no reduction.
@@
$\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)=\{a_{i},\ldots,a_{i+t\operatorname{mod}n},b_{i\operatorname{mod}n},\ldots,b_{i+t\operatorname{mod}n}\}$
@@
However, with a new technique we can still construct linear-size spanners for $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$.
For this, we need the concept of $e$-reverted edges. This construction sheds light on different ideas, such as making use of $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ to switch $A$ and $B$, that can be used for other graphs than just $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$. We start by giving the general technique and then proving that it is applicable to construct a linear-size bi-spanner for the shifted matching graph.
@@
Intuitively, for each edge $e$ there is a linear-size set of edges such that all pairs of vertices $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b^{\prime}}}\right\}$, that share an $e$-reverted edge, are connected via a short path. If ${a^{\prime}}\preceq_{b}{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)}$, the path $a^{\prime}\to b\to\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)\to b^{\prime}$ is temporal, since $b^{\prime}$ is also the latest neighbor for $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)$. If ${\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}\preceq_{a}{b^{\prime}}$, the path $a^{\prime}\to\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)\to a\to b^{\prime}$ is temporal, since $a^{\prime}$ is also the earliest neighbor $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)$. See Figure 5.
@@
Notice that, no matter which $a^{\prime}\in A$ and $b^{\prime}\in B$ we choose, the paths have many edges in common. We always use edges incident to $a$ or $b$ and $\operatorname{\pi^{\text{--}}}$ or $\operatorname{\pi^{+}}$. We use this to construct a small bi-spanner for $D$, if $\operatorname{NotRev}_{{e}}$ is small.
@@
We show the statement only for $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)=0$, the proof for label $n-1$ follows analogously. Let wlog. $a=a_{0}$ and $b=b_{0}$ and let $i,j\in\{0,\ldots,{n}-1\}$. To show that $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}$ is $e$-reverted, we distinguish two cases.
@@
Now we connect the concept of $e$-reverted edges to $c$-pivot-edges from Section 5
@@
Lemma 6.12 tells us something about where to look for edges $e$ with many $e$-reverted edges. In fact, only edges that are either among the earliest or the latest edges of the incident vertices are suitable candidates for Theorem 6.7, which are the least likely to be $c$-pivot-edges. This gives us an intuition, why the two techniques can be applied in different situations. This intuition is supported by the concrete example of shifted matching graphs.

### v2

@@ 4 Structure of Bi-Clique Instances
Using Theorem 3.1, from now on we only consider bi-cliques. Our first structural insight for this type of instance was already used by Casteigts, Peters, and Schoeters [6] on special reduced instances of bi-cliques. Remember, that we call an instance extremally matched if the set of all earliest edges $\operatorname{\pi^{\text{--}}}$ and the set of all latest edges $\operatorname{\pi^{+}}$ each form a perfect matching. We give a general reduction rule which reduces the graph in case the instance is not extremally matched. This structural property, that we can assume the existence of such perfect matchings or reduce the instance, is useful for almost all further constructions. For example, it enables us to reduce the instance size, when the number of vertices per side is not equal. Additionally, we can think of the sets $A$ and $B$ of the bi-clique as interchangeable. For example, consider any set of edges that includes temporal paths from some set $A^{\prime}\subseteq A$ to all vertices in $A$. We can extend these paths to reach all vertices in $B$ by including the set of all latest edges, because it is a perfect matching between $A$ and $B$.
@@
Note that this corresponds to the definition in [6], but transferred to the case of bi-spanners, where all vertices in $A$ only need to reach vertices in $B$. Concretely, this means that vertex $a$ can reach vertex $a^{\prime}$ via the earliest edge of $a^{\prime}$ on a temporal path of length 2. Therefore, if this configuration occurs in the bi-clique, we can remove $a$ and include the edges $\mathopen{}\mathclose{{}\left\{{a},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}}\right\}$ and $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}}\right\}$ into our bi-spanner. In the final solution for the whole bi-clique, vertex $a$ then has a temporal path to every vertex $b\in B$ via vertex $a^{\prime}$. Similarly, if a vertex $b\in B$ is dismountable, every temporal path that ends in $b^{\prime}$ can be extended to be a temporal path ending in vertex $b$, and we can remove $b$ from the bi-clique.
@@
Since from now on, both parts of $D$ have the same size after removing dismountable vertices, we write $n\coloneqq|A|=|B|$ and call $n$ the size of the instance. This justifies also that the definition of $\operatorname{\mathcal{D}}$ assumes equal part sizes and gives the following corollary.
@@ 5 Pivotable Bi-Cliques
However, there are instances without this structure [6]. Thus, we study an even more widely applicable structure, based on a what we call a $c$-pivot-edge or partial pivot-edge (when omitting the concrete parameter). We exploit partial pivot-edges to apply the divide-and-conquer paradigm to find linear-size temporal spanners. To the best of our knowledge, this paradigm has not yet been applied to this problem. As we see later, the absence of a partial pivot-edge yields strong structural properties and it heavily restricts the class of temporal cliques where no linear spanner can be found using current techniques.
5.1 Partial Pivot-Edges
@@
Let $e=\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}\in{A}\otimes{B}$. All vertices in $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ can be organized in a tree $T$ rooted at $a$, s.t. $T$ only has edges with label at least $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$ and for each $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ there are temporal paths $a\leadsto v$ and $b\leadsto v$ in $T$. The analogous result holds for vertices in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$.
The proof of Lemma 5.3 is already known [19] and it relies on constructing a foremost tree for $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ (and its inverted version for $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right))$. It was considered in [16] and lately has been used in [7], where foremost trees are applied to temporal cliques. With this idea, we can temporally connect all $u\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ to all $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ using a linear number of edges.
@@
As $32\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+2\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-3\leq 8n-$ 3, we see that only a linear number of edges is included to split the instance. With this, we can show that if for some edge $e\in{A}\otimes{B}$ the sets $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ and $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ overlap, we can reduce the instance significantly. We name the set ${\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}$ the pivot-set of $e$. If $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq 2cn$ for $c\in(0,1]$, we call $e$ a $c$-pivot-edge. When omitting the concrete constant $c$, we refer to $e$ as a partial pivot-edge.
Theorem 5.8.
Let $c\in(0,1]$. Let $D=(A,B,\lambda)$ be an extremally matched bi-clique with $n\coloneqq|A|=|B|$. If there is a $c$-pivot-edge in $D$, we can split the instance into two sub-instances with overall number of vertices reduced by $2cn$. Per removed vertex, we include at most $\frac{4}{c}$ edges. If applicable recursively, we get a spanner of size $\frac{8}{c}n$.
@@
Let $e\in{A}\otimes{B}$ be such that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq 2cn$. Note that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|=\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cup\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq\mathopen{}\mathclose{{}\left(2+2c}\right)n$. By Theorem 5.6, the sub-instances have sides of overall size at most $2n-\frac{2+2c}{2}n=\mathopen{}\mathclose{{}\left(1-c}\right)n$. This means that $2cn$ vertices have been removed using at most $38n-$ 3 edges overall. Per such vertex, we include $\frac{8n-3}{2cn}\leq\frac{4}{c}$ edges.
To recursively apply Theorem 5.6, observe that we can reduce any dismountable vertex until both sub-instances are extremally matched. For each vertex removed in this way, we include $2\leq\frac{4}{c}$ edges. As the number of edges included per removed vertex in both ways does not depend on $n$, we include at most $\frac{4}{c}\mathopen{}\mathclose{{}\left|A\sqcup B}\right|=\frac{8}{c}n$ edges.
@@
5.2 Graph Properties leading to Partial Pivot-Edges
Partial pivot-edges generalize pivotable vertices and yield a number of new, interesting structural properties in residual instances. In this section, we examine structures that, if present, imply that there is at least one partial pivot-edge.
We first look at one particular case of partial pivot-edges, which we call steep edges. Intuitively, an edge is steep, if it belongs to the later edges for one of the incident vertices and to the earlier edges for the other one.
@@ 5.2 Graph Properties leading to Partial Pivot-Edges
While steep edges demonstrate how an intuitively common structure cannot be there if assume there are no partial pivot-edges. However, in randomly generated instances we expect edges not to be steep. Assume we add edges randomly to the graph in the order of their labels. Then, when adding the edge $\mathopen{}\mathclose{{}\left\{v,w}\right\}$, both $v$ and $w$ should already have a similar number of edges. In fact, one can prove that with high probability, there is no $c$-steep in the graph.
Yet, as every $c$-steep edge is a $c$-pivot-edge, we want to analyze the graphs without such edges in greater detail. We first look at more simple properties showing that any node only has a limited number of neighbors for which it is at least the $i$-th and at most the $j$-th neighbor.
@@ 6 Bi-Cliques With Reverted Edges
We introduced partial pivot-edges and observed some interesting structural properties in bi-cliques which contain no suitable $c$-pivot-edge. The first hope would be that every bi-cliques admits such a $c$-pivot-edge. Unfortunately, this is not the case and we provide a graph class, the shifted matching graph, in which for every edge, the pivot-set is as small as possible (i.e., $\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)=\{a,b\}$) and none of our reduction rules from Section 4 apply.
Still, we present a novel technique, called $e$-reverted edges, to construct small spanners for shifted matching graphs which also provides solutions for a significantly larger class of bi-cliques. We relate the technique to partial pivot-edges from Section 5 and argue why $e$-reverted edges solve a class distinct from partial pivot-edges, beyond the concrete example of shifted matching graphs. Unfortunately, also this is not applicable to all graph classes, which we later demonstrate with a class that we call product graphs in Section 7.
@@
Figure 4 shows the shifted matching graph for $n=4$ vertices on both sides. In shifted matching graphs, the earliest and latest neighbors form a matching, thus we cannot apply the reduction from Theorem 4.7. Note that for all $i\in\{0,\ldots,{n}-1\}$, we have $\pi^{-}(a_{i})=b_{i}$ and $\pi^{+}(a_{i})=b_{i-1\operatorname{mod}n}$. Additionally, we observe that each edge at time $t$ has the form $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{i+t\operatorname{mod}n}}}\right\}$. Thus, the edges with label $t$ form a perfect matching and $a_{i}$ sees the $b_{j}$ in ascending order of indices, circularly shifted by $i$ places compared to $a_{0}$.
@@
We claim that this graph has no partial pivot-edges that are suitable for reduction in the sense of Theorem 5.8. To prove that this method fails on the shifted matching graph, we first characterize the set of vertices which can be reached from a vertex $a_{i}$ before and after a given time. Finally, we prove that the intersection of $\operatorname{In}$ and $\operatorname{Out}$ for all edges at $a_{i}$ is trivial.
@@
Using Lemma 6.2, we can see that using partial pivot-edges in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ leads to no reduction.
@@
$\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)=\{a_{i},\ldots,a_{i+t\operatorname{mod}n},b_{i},\ldots,b_{i+t\operatorname{mod}n}\}$
@@
However, with a new technique we can still construct linear-size spanners for $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$. For this, we need the concept of $e$-reverted edges. This construction sheds light on different ideas, such as making use of $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ to switch $A$ and $B$, that can be used for other graphs than just $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$. We start by giving the general technique and then proving that it is applicable to construct a linear-size bi-spanner for the shifted matching graph.
@@
Intuitively, for each edge $e$ there is a linear-size set of edges that connects all pairs of vertices $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b^{\prime}}}\right\}$ that form an $e$-reverted edge. If ${a^{\prime}}\preceq_{b}{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)}$, the path $a^{\prime}\to b\to\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)\to b^{\prime}$ is temporal, since $b^{\prime}$ is also the latest neighbor for $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)$. If ${\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}\preceq_{a}{b^{\prime}}$, the path $a^{\prime}\to\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)\to a\to b^{\prime}$ is temporal, since $a^{\prime}$ is also the earliest neighbor $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)$. See Figure 5.
@@
Notice that, no matter which $a^{\prime}\in A$ and $b^{\prime}\in B$ we choose, the paths only use edges incident to $a$ or $b$ and edges from $\operatorname{\pi^{\text{--}}}$ or $\operatorname{\pi^{+}}$. We use this to construct a small bi-spanner for $D$, if $\operatorname{NotRev}_{{e}}$ is small.
@@
We show the statement only for $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)=0$, the proof for label $n-1$ follows analogously. Without loss of generality, let $a=a_{0}$ and $b=b_{0}$ and let $i,j\in\{0,\ldots,{n}-1\}$. To show that $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}$ is $e$-reverted, we distinguish two cases.
@@
Now we connect the concept of $e$-reverted edges to partial pivot-edges from Section 5.
@@
Lemma 6.12 tells us something about where to look for edges $e$ with many $e$-reverted edges. In fact, only edges that are either among the earliest or the latest edges of the incident vertices are suitable candidates for Theorem 6.7, which are the least likely to be suitable partial pivot-edges. This gives us an intuition, why the two techniques can be applied in different situations. This intuition is supported by the example of shifted matching graphs.